OpenStudy (anonymous):

Solve for x: |4x + 12| = 16 x = 7, x = –7 x = 1, x = –1 x = 1, x = –7 x = –1, x = 7

OpenStudy (idealist10):

abs(4x+12)=16 4x+12=16 4x=16-12 4x=4 x=4/4=1 --------------- 4x+12=-16 4x=-16-12 4x=-28 x=-28/4 x=-7 Answer: x=1, -7. The answer is C.

OpenStudy (texaschic101):

| 4x + 12 | = 16 4x + 12 = 16 -(4x + 12) = 16 4x = 16 - 12 -4x - 12 = 16 4x = 4 -4x = 16 + 12 x = 1 -4x = 28 x = -7 I agree...answer is C

OpenStudy (anonymous):

Oh. I did it wrong then lol Thank you! Can you help with a few more?

OpenStudy (texaschic101):

possibly....can't make no promises I will know the answer

OpenStudy (anonymous):

Solve: negative 1 over 2 x + 2 = –x + 7

OpenStudy (anonymous):

and also this one: Lana makes $8 an hour plus $12 an hour for every hour of overtime. Overtime hours are any hours more than 40 hours for the week. Part A: Create an equation that shows the amount of wages earned, W, for working x hours in a week when there is no overtime. (3 points) Part B: Create an equation that shows the amount of wages earned, S, for working y hours of overtime. Hint: Remember to include in the equation the amount earned from working 40 hours. (3 points) Part C: Lana earned $368 in 1 week. How many hours (regular plus overtime) did she work? (4 points) (I am so confused and I'm also sick on top of this :( )

OpenStudy (texaschic101):

-1/2x + 2 = -x + 7 --- add x to both sides -1/2x + x + 2 = 7 -- subtract 2 from both sides -1/2x + x = 7 - 2 -- combine like terms -1/2x + 2/2x = 5 1/2x = 5 x = 5/(1/2) x = 5 * 2/1 x = 10

OpenStudy (texaschic101):

A. W = 8x B. I am not sure how to write this one... C. 40 * 8 = 320 regular pay 368 - 320 = 48....48/12 = 4 she worked 40 regular hrs and 4 overtime hrs

OpenStudy (texaschic101):

Hope you get to feeling better...sorry I couldn't help more